There are 7,200 identical firms in the widget industry each of which has a fixed cost that is avoidable in the long run, F=1,500, and their technology is given by q=K^(1/5)L^(3/10). Initially the input prices are w=9 and v=6. The market demand they face is QD(P)=176,030−50 P. Find the long run equilibrium.
Answer #1

Date 1 0 = ()**340 w= 19 0=6 C = wltok =9L+6k MPL = 3 (9s ()*70 to < (6Y5 (2%0 Mpx = 1 (K) 46 L% MP2 ()s (4)%10 MPK CS Scan= 3x5 K lo L MPL 3k To minimize the cost following condition must be satisfied 1 을슨응 sealthCamScannermDote 0 = * (2140 © = (_)** (13% [= 02] K= 02 C = Wht rk C = 9 @²+682 C= 15 9² Total cost = Fixed cast & Vanable cast T.C 3.15To determine long run equilibrium we must find daw point of fiems alwage cast curree The min occurs where AC=MC 1500 +15° = 36 d = 176030-50P od = 65 176030-50P= 72000 50p=104030 IP=2080.6 CSScanned with CamScanner